### v1

# Arithmetic properties and zeros of the Bergman kernel on a class of quotient domains

Abstract.
An effective formula for the Bergman kernel on ${\mathbb{H}}_{\gamma}=\{|z_{1}|^{\gamma}<|z_{2}|<1\}$ is obtained for rational $\gamma=\frac{m}{n}>1$. The formula depends on arithmetic properties of $\gamma$, which uncovers new symmetries and clarifies previous results. The formulas are then used to study the Lu Qi-Keng problem. We produce sequences of rationals $\gamma_{j}\searrow 1$, where each ${\mathbb{H}}_{\gamma_{j}}$ has a Bergman kernel with zeros (while ${\mathbb{H}}_{1}$ is known to have a zero-free kernel), resolving an open question on this domain class.
The first author was supported in part by Austrian Science Fund (FWF) grants: DOI 10.55776/I4557 and DOI 10.55776/P36884.
2020 Mathematics Subject Classification: 32A25 (Primary); 32A60, 30C15 (Secondary)
1. Introduction
Given a domain (an open connected set) $\Omega\subset{\mathbb{C}}^{n}$, the Bergman space $A^{2}(\Omega)$ is the subspace of $L^{2}(\Omega)$ consisting of holomorphic functions. The Bergman space is itself a Hilbert space, so there is a canonical orthogonal projector $P_{\Omega}:L^{2}(\Omega)\to A^{2}(\Omega)$, called the Bergman projection. This projection is carried out by integrating against the Bergman kernel, the function $K_{\Omega}:\Omega\times\Omega\to{\mathbb{C}}$ serving as the reproducing kernel of $A^{2}(\Omega)$:
[…]
For fixed $w\in\Omega$, the function $K_{\Omega}(\cdot,w)\in A^{2}(\Omega)$; it is conjugate symmetric (i.e. $K_{\Omega}(z,w)=\overline{K_{\Omega}(w,z)}$) and satisfies many additional extremal and invariance properties; see [20].
For $\gamma>0$, define a family of bounded, non-smooth, weakly pseudoconvex domains by
${\mathbb{H}}_{\gamma}=\{(z_{1},z_{2})\in{\mathbb{C}}^{2}:|z_{1}|^{\gamma}<|z_{2}|<1\}.$ (1.2)
Bergman theory on these domains has been studied extensively in recent years; see [9, 13, 14, 15, 7, 8, 3]. When $\gamma$ is rational, ${\mathbb{H}}_{\gamma}$ can be realized as a quotient of ${\mathbb{D}}\times{\mathbb{D}}^{*}$ induced by a monomial mapping; see Section 1.2. The purpose of this paper is two fold: (1) to give an effective formula for Bergman kernel of ${\mathbb{H}}_{\gamma}$ for every rational $\gamma>1$, and (2) to use the formulas to exhibit new symmetries and locate zeros of the kernel in ${\mathbb{H}}_{\gamma}\times{\mathbb{H}}_{\gamma}$.
The question about the existence of zeros of Bergman kernel (the Lu Qi-Keng problem) goes back many decades and was originally motivated by the global well-definedness of Bergman representative coordinates; see [4, 22]. While simple domains such as the ball and polydisc have nowhere vanishing Bergman kernels, an interesting open problem is to ask when the taking of quotients of such domains induces zeros in the quotient Bergman kernels. This has been investigated in special cases (e.g., certain ball quotients; symmetrized polydiscs; the tetrablock and its variants; [5, 16, 23, 26, 17]) but a unified approach to this problem is only in its early stages.
Until the writing of this paper, the Lu Qi-Keng problem on the ${\mathbb{H}}_{\gamma}$ stood as follows:
  • $(a)$
    The Bergman kernel on ${\mathbb{H}}_{1}$ (the classical Hartogs triangle) is non-vanishing.
  • $(b)$
    The Bergman kernel on ${\mathbb{H}}_{\gamma}$ has zeros for $\gamma\in[2,\infty)$; see [12, 14]; also Section 1.3.
  • $(c)$
    It was unknown whether the Bergman kernel of ${\mathbb{H}}_{\gamma}$ has zeros for $\gamma\in(1,2)$.
The problem also remains only partially solved for $\gamma\in(0,1)$. However, given $\gamma$ in this range, there is a unique $\widetilde{\gamma}\in[1,\infty)$ such that ${\mathbb{H}}_{\gamma}$ is biholomorphic to ${\mathbb{H}}_{\widetilde{\gamma}}$; see [12, 14, 27]. Now since the transformation law in (2.2) shows that the non-vanishing of the Bergman kernel is preserved under biholomorphism, answering the Lu Qi-Keng problem for $\gamma\in(1,2)$ would in fact resolve the problem for all $\gamma>0$. This fact was observed in [14], after Edholm and McNeal found zeros of ${\mathbb{H}}_{\gamma}$ for $\gamma\geq 2$, but noted their approach failed to address $21<\gamma<$ 2 (see Section 1.3 below for more details). The question has remained open since.
  • (Q1)
    For which $\gamma\in(1,2)$ does the Bergman kernel have zeros?
Definition (1.2) shows ${\mathbb{H}}_{\gamma}$ varies continuously with $\gamma$, and Ramadanov’s theorem (see [21]) shows that if $\gamma_{j}\searrow 1$, the Bergman kernels $K_{\gamma_{j}}(\cdot,\cdot)$ tend to the Bergman kernel $K_{1}(\cdot,\cdot)$ of ${\mathbb{H}}_{1}$ locally normally on ${\mathbb{H}}_{1}\times{\mathbb{H}}_{1}$. Following the work of [11, 14], a weaker version of (Q1) has also been raised. Since $K_{1}$ is known to be non-vanishing,
  • (Q2)
    Is there some $\epsilon>0$ so that $K_{\gamma}(\cdot,\cdot)$ is non-vanishing on ${\mathbb{H}}_{\gamma}\times{\mathbb{H}}_{\gamma}$ for $\gamma\in(1,1+\epsilon)$?
This paper addresses both questions. We exhibit infinitely many rational $\gamma$ in $(1,2)$ so that $K_{\gamma}(\cdot,\cdot)$ has zeros inside ${\mathbb{H}}_{\gamma}\times{\mathbb{H}}_{\gamma}$ (see Section 4), marking the first progress to be made on (Q1). These $\gamma$ values accumulate at 1, answering (Q2) by proving no such $\epsilon$ can exist.
1.1. An effective kernel formula
While the existence of the Bergman kernel is guaranteed by abstract Hilbert space theory, concrete information about its properties as a function $K_{\Omega}:\Omega\times\Omega\to{\mathbb{C}}$ (asymptotics near the boundary; the existence of zeros inside $\Omega\times\Omega$ and properties of the zero set) requires hard analysis.
Our first main result is a new and effective formula for the Bergman kernel of ${\mathbb{H}}_{m/n}$, valid for any rational $\frac{m}{n}>1$. Previous work (see [14, 8, 1]) has expressed these kernels (which have been known since [14] to be rational functions) in one form or another, but a complete and effective formula for the numerator polynomial, from which it is possible to study the zero set of the kernel when $\gamma\in(1,2)$, has until now remained elusive. These past results, along with how Theorem 1.3 clarifies and improves them, are presented in Section 1.3.
Theorem 1.3.
Let $m,n\in{\mathbb{Z}}^{+}$ with $\gcd(m,n)=1$ and $m>n$. Given points $z=(z_{1},z_{2})$ and $w=(w_{1},w_{2})$ in ${\mathbb{H}}_{m/n}$, let $s=z_{1}\overline{w}_{1}$, $t=z_{2}\overline{w}_{2}$. The Bergman kernel of ${\mathbb{H}}_{m/n}$ is
$K_{m/n}(z,w)=\frac{P_{m,n}(s,t)}{m\pi^{2}(1-t)^{2}(t^{n}-s^{m})^{2}},$ (1.4a)
where the numerator is a polynomial of degree $12m-$ 1 given explicitly as
$P_{m,n}(s,t)=p_{0}(s,t)+p_{1}(s,t)+p_{2}(s,t)+p_{3}(s,t)+p_{4}(s,t),$ (1.4b)
where
$p_{0}(s,t)$ $=m^{2}s^{m-1}t^{n},$ (1.4c)
$p_{1}(s,t)$ $=\sum_{j=0}^{m-2}(j+1)(\kappa(j)+1)s^{j}t^{2n-L(j)},$ (1.4d)
$p_{2}(s,t)$ $=\sum_{j=0}^{m-2}(j+1)(m-\kappa(j)-1)s^{j}t^{2n+1-L(j)},$ (1.4e)
$p_{3}(s,t)$ $=\sum_{j=0}^{m-2}(m-j-1)(\kappa(j)+1)s^{j+m}t^{n-L(j)},$ (1.4f)
$p_{4}(s,t)$ $=\sum_{j=0}^{m-2}(m-j-1)(m-\kappa(j)-1)s^{j+m}t^{n+1-L(j)},$ (1.4g)
and the expressions $L$ and $\kappa$ are given by
$L(j)=\left\lceil\frac{1+n(j+1)}{m}\right\rceil,\qquad\kappa(j)=m+n-1+nj-mL(j).$ (1.4h)
This theorem is proved in Section 3 after employing modular arithmetic to solve several linear Diophantine equations depending on $m$ and $n$ which determine the complete list of monomials appearing in the numerator polynomial (Proposition 3.4). This information is then combined with functional identities satisfied by the quantities $L$ and $\kappa$ (Lemma 3.13) to produce the Bergman kernel formula.
The new kernel formulas allow for effective analysis in part because they improve (in fact optimize) computational efficiency of previous formulas. For example, the explicit formula for $K_{m/n}(z,w)$ given in [1, Proposition 2] (see Theorem 1.11 below) requires the computation of $(2m-1)(2n+1)$ polynomial coefficients, while Theorem 1.3 shows that all but exactly $34m-$ 3 of these coefficients are equal to zero; see Section 1.3 for more.
1.2. Monomial maps; symmetries and zeros of the Bergman kernel
We recall the point of view taken in [12, 3, 8, 1] which realizes ${\mathbb{H}}_{m/n}$ as quotients. Suppose $m,n\in{\mathbb{Z}}^{+}$ are relatively prime. The image of ${\mathbb{D}}\times{\mathbb{D}}^{*}$ under the monomial map $\Phi_{m,n}(z_{1},z_{2})=(z_{1}z_{2}^{n},z_{2}^{m})$ can be identified with a rational Hartogs triangle:
$\Phi_{m,n}({\mathbb{D}}\times{\mathbb{D}}^{*})={\mathbb{H}}_{m/n}.$
More precisely, $\Phi_{m,n}:{\mathbb{D}}\times{\mathbb{D}}^{*}\to{\mathbb{H}}_{m/n}$ is an $m$-sheeted proper holomorphic branched cover. In the language of [3, 6], $\Phi_{m,n}$ is a proper holomorphic map of quotient type with respect to a group $\Gamma_{m,n}\subset\mathrm{Aut}({\mathbb{D}}\times{\mathbb{D}}^{*})$, the deck transformation group of $\Phi_{m,n}$, which in this case is cyclic of order $m$. In light of these considerations, Theorem 1.3 demonstrates that an intricate Bergman theory can arise on quotients, even when the starting domains and maps are simple. In this case $\Phi_{m,n}$ forces arithmetic properties of the rational index $\gamma=\frac{m}{n}$ into the Bergman kernel via Bell’s transformation rule for proper maps [2].
1.2.1. Palindromic polynomials
The numerator polynomial of the Bergman kernel of ${\mathbb{H}}_{m/n}$ satisfies an interesting symmetry that assists the study of the Lu Qi-Keng problem.
Given points $z,w\in{\mathbb{H}}_{\gamma}$, we follow the notation of Theorem 1.3 and write $s=z_{1}\overline{w}_{1}$, $t=z_{2}\overline{w}_{2}$. It is easy to check via Corollary 2.8 and Proposition 2.13 that the point $(s,t)\in{\mathbb{H}}_{\gamma}$, and that every point of ${\mathbb{H}}_{\gamma}$ can be realized in this way. Now let $(s,t)$ denote coordinates on ${\mathbb{C}}^{2}$, and consider the complex line $\{s=t\}$. This line intersects the interior of ${\mathbb{H}}_{\gamma}=\{|s|^{\gamma}<|t|<1\}$ for any $\gamma>1$, where the intersection is a punctured disc $\Delta^{*}$ lying obliquely inside of ${\mathbb{H}}_{\gamma}$ (see Figure 1):
$\Delta^{*}=\{(s,s)\in{\mathbb{H}}_{\gamma}:s\in{\mathbb{D}}^{*}\}.$ (1.5)
Now restrict the numerator of (1.4a) to the line $\{s=t\}$. The one variable polynomial $s\mapsto P_{m,n}(s,s)$ has a root of order $12n-$ 1 at the origin, suggesting the following definition
$Q_{m,n}(s):=s^{1-2n}P_{m,n}(s,s).$ (1.6)
Thus, roots of the one variable polynomial $Q_{m,n}(s)$ inside ${\mathbb{D}}^{*}\subset{\mathbb{C}}$ are in one-to-one correspondence with the roots of $P_{m,n}(s,t)$ restricted to $\Delta^{*}\in{\mathbb{H}}_{\gamma}$. Understanding these roots is made easier due to the following unexpected symmetry in the coefficients of $Q_{m,n}$:
Theorem 1.7.
Given relatively prime integers $m>n$, $Q_{m,n}$ is a palindromic polynomial of degree $2m-2n$, i.e.,
$Q_{m,n}(s)=s^{2m-2n}Q_{m,n}(s^{-1}).$ (1.8)
Theorem 1.7 is obtained as part of Proposition 3.21, which establishes identities analogous to (1.8) on pieces of the full polynomial. The palindromic property of $Q_{m,n}$ allows for indirect study of roots inside ${\mathbb{D}}^{*}$ by examining the polynomial’s behavior on the unit circle.
This circle of ideas pays off in Section 4 where it is shown that $Q_{m,n}$ is non-vanishing on the unit circle for all relatively prime pairs $m>n$ such that $m-n=1$ or $2$:
Theorem 1.9.
If $m-n=1$ or $2$, then $Q_{m,n}$ has exactly half of its roots inside the unit disc. Consequently, the Bergman kernel of ${\mathbb{H}}_{m/n}$ has zeros inside ${\mathbb{H}}_{m/n}\times{\mathbb{H}}_{m/n}$.
Proofs of different pieces of this theorem are found in Section 4.1 and 4.2. As a consequence of this result, we obtain infinitely many rationals accumulating at $\gamma=1$ such that the Bergman kernel has zeros, resolving (Q2) above with a negative answer.
$|t|$$|s|$$\{s=t\}\,\,$
Figure 1. The punctured disc $\Delta^{*}$ lying obliquely inside ${\mathbb{H}}_{\gamma}$.
1.3. Previous results
The Bergman kernel of ${\mathbb{H}}_{\gamma}$ was first computed for $\gamma\in{\mathbb{Z}}^{+}$ by Edholm in [12], then for $\gamma=\frac{m}{n}\in{\mathbb{Q}}^{+}$ by Edholm and McNeal in [14], where it was expressed as a sum of $m$ subkernels, all rational functions with a common denominator. Estimates on these subkernels imply a sharp bound on the full Bergman kernel, which then leads to a characterization of $L^{p}$-mapping properties of the Bergman projection based on the rationality/irrationality of $\gamma$; see [14, Theorem 1.1 and 1.2].
Because the $L^{p}$ results were deducible directly from the kernel estimates, little attempt was made to further analyze the numerator of the Bergman kernel, or re-write it in a way that could make symmetries like Theorem 1.7 easier to detect. The only result in this direction was [14, Theorem 3.3], where it was noted that all but one of the subkernels vanish identically on the variety $\{s=0\}$, meaning the full Bergman kernel restricted to this variety is given by a particularly simple formula:
$K_{\gamma}((0,z_{2}),(0,w_{2}))=\frac{1+(\gamma-1)z_{2}\overline{w}_{2}}{\gamma\pi^{2}z_{2}\overline{w}_{2}(1-z_{2}\overline{w}_{2})^{2}}.$
(This formula is valid for all $\gamma>0$, including irrationals.) When $\gamma>2$, it is possible by direct inspection to find $z_{2},w_{2}$ with $10<|z_{2}|,|w_{2}|<$ 1 for which the Bergman kernel vanishes. But for $0<\gamma\leq 2$, it is not possible to find such $z_{2},w_{2}$, meaning that those Bergman kernels are non-vanishing on $\{s=0\}$. (The $\gamma=2$ case had previously been addressed by Edholm in [12, Corollary 4.2], where an explicit zero of $K_{2}$ inside ${\mathbb{H}}_{2}\times{\mathbb{H}}_{2}$ was produced.) This approach of looking at the kernel restricted to a variety inspired the restriction of the Bergman kernel to $\{s=t\}$ considered in the present paper.
1.3.1. The Bergman kernel in terms of the function $D_{m}$
Another approach to the Bergman kernels of ${\mathbb{H}}_{m/n}$ comes from Chakrabarti et al. [8], who obtained formulas as special cases of calculations on so-called Signature 1 domains in ${\mathbb{C}}^{n}$. In this work, proper holomorphic maps are used to cover the domains under consideration by simpler domains for which the Bergman kernel is already explicitly known. Bell’s transformation law (see [2]) can then be applied to express the kernel on the target as a linear combination of functions closely related to the the Bergman kernel on the source. The difficulty with this approach is parsing this linear combination into a digestible formula that can then be used to effectively study some other problem (e.g. mapping properties of the Bergman projection or locating zeros).
One of the main observations of [8] is that the Bergman kernels of Signature 1 domains (including the ${\mathbb{H}}_{m/n}$) are rational functions, whose numerators can be succinctly expressed using the following function defined on the integers:
$D_{m}(\beta)=\begin{cases}\beta+1,&0\leq\beta\leq m-1,\\ 2m-1-\beta,&m\leq\beta\leq 2m-2,\\ 0,&\mathrm{otherwise}.\end{cases}$ (1.10)
This function arises the coefficients of the following polynomial equation:
$\left(\frac{1-s^{m}}{1-s}\right)^{2}=\sum_{\beta=0}^{2m-2}D_{k}(\beta)s^{k}.$
We now present these Bergman kernel formulas in terms of $D_{m}$, using the description given by Almughrabi [1], who simplified the original presentation in this form given in [8]:
Proposition 1.11 ([1], Proposition 2).
Let $m,n\in{\mathbb{Z}}^{+}$ with $m>n$ and $\gcd(m,n)=1$. Letting $s=z_{1}\overline{w}_{1}$, $t=z_{2}\overline{w}_{2}$, the Bergman kernel of ${\mathbb{H}}_{m/n}$ is given by
$K_{m/n}(z,w)=\sum_{\beta\in{\mathbb{Z}}^{2}}\frac{D_{m}(\beta_{1})D_{m}(m\beta_{2}+n\beta_{1}+m+n-1-2mn)s^{\beta_{1}}t^{\beta_{2}}}{\pi^{2}m(t^{n}-s^{m})^{2}(1-t)^{2}}.$ (1.12)
In this formula, the terms in the numerator polynomial are only possibly non-zero when
$0\leq\beta_{1}\leq 2m-2,\qquad\mathrm{and}\qquad 0\leq\beta_{2}\leq 2n.$ (1.13)
While this formula expressing the numerator polynomial in terms of $D_{m}$ provides some insight into the algebra of the covering maps $\Phi_{m,n}:{\mathbb{D}}\times{\mathbb{D}}^{*}\to{\mathbb{H}}_{m,n}$, basic information about the polynomial: the degree, which monomials appear in the sum with non-zero coefficients, or what these coefficients are, remains somewhat obfuscated. The kernel formulas given in Theorem 1.3 resolve these questions entirely. The condition in (1.13) requires the computation of $(2m-1)(2n+1)$ polynomial coefficients, but formuals (1.4c) - (1.4g), see that only $34m-$ 3 of these coefficients are non-zero. For large $n$, this represents an enormous improvement in the efficiency of the formula presentation.
2. Background
2.1. Reinhardt domains
A domain $\Omega\subset{\mathbb{C}}^{n}$ is called Reinhardt (with respect to the origin) when it admits a natural $n$-torus action allowing for independent rotation in each coordinate. In other words,
$(z_{1},z_{2},\cdots,z_{n})\in\Omega\quad\Longleftrightarrow\quad(e^{i\theta_{1}}z_{1},e^{i\theta_{2}}z_{2},\cdots,e^{i\theta_{n}}z_{n})\in\Omega$
for any choice of $\theta_{1},\dots,\theta_{n}\in{\mathbb{R}}$. In ${\mathbb{C}}^{1}$ the only Reinhardt domains are (1) discs, (2) the full plane, (3) annuli, (4) the plane with a closed disc removed, (5) once punctured discs with the center removed, and (6) the once punctured plane. In each of these settings, holomorphic functions are represented by globally valid Laurent expansions, but every such domain family has its own distinct flavor of holomorphic function theory. This distinction can be drastic, as the contrast between the classical theories of entire functions and holomorphic functions on the unit disc indicates.
Holomorphic functions on Reinhardt domains in ${\mathbb{C}}^{n}$ also admit globally valid Laurent expansions, and as before, two distinct Reinhardt domains often give rise to substantially different holomorphic function theories; e.g., the function theory on the unit ball [25] versus the unit polydisc [24]. Properties of the Bergman kernel (boundary asymptotics; the existence of zeros and shape of the zero set) and Bergman projection (its behavior acting in $L^{p}$, Sobolev, Hölder classes) give a way to gauge this.
A domain is said to be pseudoconvex when it is the domain of existence for some holomorphic function that cannot be analytically extended to a larger domain. Pseudoconvexity is a fundamental notion in several complex variables, and it has several equivalent geometric characterizations; see [20]. In the Reinhardt setting it is particularly easy to characterize:
Given a Reinhardt domain $\Omega\subset{\mathbb{C}}^{n}$, define its logarithmic shadow to be the set
$\mathrm{Log}(\Omega):=\{(\log|z_{1}|,\cdots,\log|z_{n}|)\in{\mathbb{R}}^{n}:z\in\Omega\,\,\mathrm{with\,\,each}\,\,z_{j}\neq 0\}.$
A domain is called weakly relatively complete if, for each $j$ with $\Omega\cap\{z_{j}=0\}\neq\varnothing$, that $(z^{\prime},0,z^{\prime\prime})\in\Omega$ whenever there exists some $(z^{\prime},z_{j},z^{\prime\prime})\in\Omega\subset{\mathbb{C}}^{j-1}\times{\mathbb{C}}\times{\mathbb{C}}^{n-j}$.
Proposition 2.1 ([19], Proposition 1.11.13).
A Reinhardt domain $\Omega\subset{\mathbb{C}}^{n}$ is pseudoconvex if and only if $\mathrm{Log}(\Omega)\subset{\mathbb{R}}^{n}$ is a convex set and $\Omega$ is weakly relatively complete.
It is easily verified that the domains ${\mathbb{H}}_{\gamma}$ are pseudoconvex.
2.2. Bergman kernel properties
Let $U$ and $V$ be two domains in ${\mathbb{C}}^{n}$ and suppose there is a biholomorphic map $\Phi:U\to V$. The Bergman kernels satisfy the well known relation
$K_{U}(z,w)=\det\Phi^{\prime}(z)\cdot K_{V}(\Phi(z),\Phi(w))\cdot\overline{\det\Phi^{\prime}(w)}.$ (2.2)
Since $\det\Phi^{\prime}$ is necessarily non-vanishing for $\Phi$ a biholomorphism (see [20]) we see that $K_{U}$ has zeros inside $U\times U$ if and only if $K_{V}$ has zeros inside $V\times V$.
Since the map $(z_{1},z_{2})\mapsto\big{(}\frac{z_{1}}{z_{2}},z_{2}\big{)}$ gives a biholomorphism from ${\mathbb{H}}_{1}$ to ${\mathbb{D}}\times{\mathbb{D}}^{*}$ (a domain with a non-vanishing Bergman kernel), we can immediately conclude $K_{{\mathbb{H}}_{1}}(z,w)$ is non-vanishing inside ${\mathbb{H}}_{1}\times{\mathbb{H}}_{1}$. Alternatively, this can be seen directly from the formula
$K_{{\mathbb{H}}_{1}}(z,w)=\frac{z_{2}\overline{w}_{2}}{\pi^{2}(1-z_{2}\overline{w}_{2})^{2}(z_{2}\overline{w}_{2}-z_{1}\overline{w}_{1})^{2}}.$
Given a Reinhardt domain $\Omega\subset{\mathbb{C}}^{n}$ and a multi-index $\alpha\in{\mathbb{Z}}^{n}$ with corresponding monomial $e_{\alpha}(z)=z^{\alpha}$, we define the set of $L^{2}$-allowable indices to be
${\mathcal{S}}_{2}(\Omega):=\{\alpha\in{\mathbb{Z}}^{n}:e_{\alpha}\in A^{2}(\Omega)\}.$
Since the square-integrable monomials form a basis for the Bergman space, we are able to write the Bergman kernel as
$K_{\Omega}(z,w)=\sum_{\alpha\in{\mathcal{S}}_{2}(\Omega)}\frac{e_{\alpha}(z)\overline{e_{\alpha}(w)}}{\left\|e_{\alpha}\right\|_{2}^{2}}=\sum_{\alpha\in{\mathcal{S}}_{2}(\Omega)}\frac{(z_{1}\overline{w}_{1})^{\alpha_{1}}\cdots(z_{n}\overline{w}_{n})^{\alpha_{n}}}{\left\|e_{\alpha}\right\|_{2}^{2}},$ (2.3)
where the series converges locally normally on $\Omega\times\Omega$. Note that on Reinhardt domains like ${\mathbb{H}}_{\gamma}$ – where the boundary intersects the center of Reinhardt symmetry – certain negative powers are allowed to appear on square-integrable monomials. But any perturbation of $\gamma$ will cause the set ${\mathcal{S}}_{2}({\mathbb{H}}_{\gamma})\subset{\mathbb{Z}}^{2}$ to either lose or gain infinitely many indices, a fact that underlies the drastic difference in Bergman kernels of ${\mathbb{H}}_{\gamma_{1}}\neq{\mathbb{H}}_{\gamma_{2}}$, even for $|\gamma_{1}-\gamma_{2}|$ arbitrarily small.
Equation (2.3) shows the Bergman kernel on a Reinhardt domain can be written as a function of the variables $t_{j}=z_{j}\overline{w}_{j}$. In other words, if $\psi:{\mathbb{C}}^{n}\times{\mathbb{C}}^{n}\to{\mathbb{C}}^{n}$ is the map
$\psi(z,w)=(z_{1}\overline{w}_{1},\cdots,z_{n}\overline{w}_{n}),$
the Bergman kernel factors as $K_{\Omega}=k_{\Omega}\circ\psi$, where $k_{\Omega}$ is a function defined on the closely related Reinhardt domain
$\widetilde{\Omega}=\{t\in{\mathbb{C}}^{n}:\exists\,z,w\in\Omega\,\,\mathrm{with}\,\,t=(z_{1}\overline{w}_{1},\cdots,z_{n}\overline{w}_{n})\}.$ (2.4)
When $\Omega$ is pseudoconvex, $\widetilde{\Omega}$ yields a simple description. Following [6], for $m>0$ we define the $m$-th Reinhardt power of $\Omega$ to be
$\Omega^{(m)}:=\left\{z\in{\mathbb{C}}^{n}:(|z_{1}|^{\frac{1}{m}},\cdots,|z_{n}|^{\frac{1}{m}})\in\Omega\right\}.$ (2.5)
Proposition 2.6.
Let $\Omega\subset{\mathbb{C}}^{n}$ be a pseudoconvex Reinhardt domain. Then $\widetilde{\Omega}=\Omega^{(2)}$.
Proof.
The containment $\Omega^{(2)}\subset\widetilde{\Omega}$ always holds. Indeed, given $\zeta\in\Omega^{(2)}$, we have
$\zeta:=(r_{1}e^{i\theta_{1}},\dots,r_{n}e^{i\theta_{n}})\in\Omega^{(2)}\qquad\Longleftrightarrow\qquad z:=\left(\sqrt{r_{1}},\dots,\sqrt{r_{n}}\right)\in\Omega.$
Now take $w=z$ and set $t:=(z_{1}\overline{w}_{1},\cdots,z_{n}\overline{w}_{n})=(r_{1},\cdots,r_{n})$, which is contained in $\widetilde{\Omega}$ by definition (2.4). But since $\widetilde{\Omega}$ is Reinhardt, $\zeta\in\widetilde{\Omega}$.
The containment $\widetilde{\Omega}\subset\Omega^{(2)}$ requires pseudoconvexity. Let $t=(t_{1},\dots,t_{n})\in\widetilde{\Omega}$ and choose
$z=(z_{1},\dots,z_{n})\in\Omega,\qquad w=(w_{1},\dots,w_{n})\in\Omega,$
such that $t_{j}=z_{j}\overline{w}_{j}$ for all $j$. For now, assume that each $t_{j}\neq 0$. Since $\Omega$ is pseudoconvex, the set $\mathrm{Log}(\Omega)\subset{\mathbb{R}}^{n}$ is convex and therefore the midpoint
$\left(\frac{\log|z_{1}|+\log|w_{1}|}{2},\dots,\frac{\log|z_{n}|+\log|w_{n}|}{2}\right)\in\mathrm{Log}(\Omega).$ (2.7)
Upon noting that $\mathrm{Log}\big{(}\Omega^{(2)}\big{)}=2\,\mathrm{Log}\big{(}\Omega\big{)}:=\left\{2\eta:\eta\in\mathrm{Log}(\Omega)\right\}$, we see from (2.7) that
$(\log|t_{1}|,\dots,\log|t_{n}|)=(\log|z_{1}|+\log|w_{1}|,\dots,\log|z_{n}|+\log|w_{n}|)\in\mathrm{Log}\big{(}\Omega^{(2)}\big{)},$
which shows that $t=(t_{1},\dots,t_{n})\in\Omega^{(2)}$.
When some of the components $t_{j}$ equal 00, we take $z_{j}=w_{j}=0$, and repeat the midpoint argument above on the logarithmic shadow of the lower dimensional Reinhardt domain obtained by intersecting $\Omega$ with the coordinate axes of the zeroed out components. This lower dimensional domain is also pseudoconvex and its $m$-th Reinhardt powers are obtained by intersecting $\Omega^{(m)}$ with the same coordinate axes. The rest of the argument goes through mutatis mutandis, confirming again that $t\in\Omega^{(2)}$. ∎
Proposition 2.6 has the following simple consequence
Corollary 2.8.
Given a pseudocovex Reinhart domain $\Omega\subset{\mathbb{C}}^{n}$, the map $\psi$ restricts to a surjection of $\Omega\times\Omega$ onto $\Omega^{(2)}$, and there is a holomorphic function $k_{\Omega}:\Omega^{(2)}\to{\mathbb{C}}$ so that the following diagram commutes
${\Omega\times\Omega}$${\Omega^{(2)}}$${{\mathbb{C}}}$$\scriptstyle{\psi}$$\scriptstyle{K_{\Omega}}$$\scriptstyle{k_{\Omega}}$ (2.9)
Proof.
Since $\psi(\Omega\times\Omega)=\widetilde{\Omega}=\Omega^{(2)}$, the factorization of $K_{\Omega}$ illustrated in (2.9) holds. The function $k_{\Omega}$ is holomorphic on $\Omega^{(2)}$ because the power series
$k_{\Omega}(t_{1},\cdots,t_{n})=\sum_{\alpha\in{\mathcal{S}}_{2}(\Omega)}\frac{t_{1}^{\alpha_{1}}\cdots t_{n}^{\alpha_{n}}}{\left\|e_{\alpha}\right\|_{2}^{2}}$ (2.10)
converges locally normally on $\Omega^{(2)}$ – this is a direct consequence of the local normal convergence of the Bergman kernel (2.3) on $\Omega\times\Omega$. ∎
Definition 2.11.
Given a pseudoconvex Reinhardt domain $\Omega\subset{\mathbb{C}}^{n}$, we call the function $k_{\Omega}:\Omega^{(2)}\to{\mathbb{C}}$ the holomorphic factor of the Bergman kernel.
Corollary 2.12.
Let $\Omega\subset{\mathbb{C}}^{n}$ be a pseudoconvex Reinhardt domain. The zero set of the Bergman kernel $K_{\Omega}$ in $\Omega\times\Omega$ is the pullback via $\psi$ of the zero set in $\Omega^{(2)}$ of the holomorphic factor $k_{\Omega}$. In particular, $K_{\Omega}$ has zeros in $\Omega\times\Omega$ if and only if $k_{\Omega}$ has zeros in $\Omega^{(2)}$.
Proposition 2.13.
For $m>0$, $\gamma>0$ the generalized Hartogs triangle satisfies ${\mathbb{H}}_{\gamma}^{(m)}={\mathbb{H}}_{\gamma}$, and the holomorphic factor of the Bergman kernel is itself a function $k_{\gamma}:{\mathbb{H}}_{\gamma}\to{\mathbb{C}}$.
Proof.
Given $m,\gamma>0$, we have the following chain of equivalent statements:
$z\in{\mathbb{H}}_{\gamma}^{(m)}\quad\Longleftrightarrow\quad\big{(}|z_{1}|^{\frac{1}{m}},|z_{2}|^{\frac{1}{m}}\big{)}\in{\mathbb{H}}_{\gamma}\quad$ $\Longleftrightarrow\quad|z_{1}|^{\frac{\gamma}{m}}<|z_{2}|^{\frac{1}{m}}<1$
$\Longleftrightarrow\quad|z_{1}|^{\gamma}<|z_{2}|<1\qquad\Longleftrightarrow\quad z\in{\mathbb{H}}_{\gamma}.$
Since ${\mathbb{H}}_{\gamma}^{(2)}={\mathbb{H}}_{\gamma}$, the holomorphic factor $k_{\gamma}$ is defined on the original domain. ∎
3. An effective Bergman kernel formula
The goal here is to start with the numerator polynomial given in (1.12):
$\sum_{\beta\in{\mathbb{Z}}^{2}}D_{m}(\beta_{1})D_{m}(m\beta_{2}+n\beta_{2}+m+n-1-2mn)s^{\beta_{1}}t^{\beta_{2}}:=P_{m,n}(s,t),$ (3.1)
and make it effective by eliminating the $D_{m}$ function that was defined in (1.10). To condense the presentation of (3.1), we introduce the quantities
$\kappa(\beta_{1},\beta_{2})$ $:=m\beta_{2}+n\beta_{1}+m+n-1-2mn,$ (3.2)
$C(\beta_{1},\beta_{2})$ $:=D_{m}(\beta_{1})D_{m}(\kappa(\beta_{1},\beta_{2})).$ (3.3)
3.1. Determining non-zero coefficients
We determine all pairs $(\beta_{1},\beta_{2})\in{\mathbb{Z}}^{2}$ such that $C(\beta_{1},\beta_{2})\neq 0$.
Proposition 3.4.
Fix a rational number $\frac{m}{n}>1$ with $\gcd(m,n)=1$ and let $C(\beta_{1},\beta_{2})$ be defined as in (3.3). The pairs $(\beta_{1},\beta_{2})\in{\mathbb{Z}}^{2}$ for which $C(\beta_{1},\beta_{2})\neq 0$ are obtained explicitly.
  • (0)
    $C(\beta_{1},\beta_{2})\neq 0$ if and only if both $20\leq\beta_{1}\leq 2m-$ 2 and $20\leq\kappa(\beta_{1},\beta_{2})\leq 2m-$ 2.
Case 1: $\beta_{1}=m-1$ or $\kappa(\beta_{1},\beta_{2})=m-1$.
  • (1a)
    If $\beta_{1}=m-1$, the only possible $\beta_{2}$ value for which $C(\beta_{1},\beta_{2})\neq 0$ is $\beta_{2}=n$, in which case $\kappa(\beta_{1},\beta_{2})=m-1$.
  • (1b)
    Suppose that $\beta_{1},\beta_{2}$ are chosen so that $C(\beta_{1},\beta_{2})\neq 0$ and $\kappa(\beta_{1},\beta_{2})=m-1$. Then $\beta_{1}=m-1$ and $\beta_{2}=n$.
For what follows, define
$L(j)=\left\lceil\frac{1+n(j+1)}{m}\right\rceil.$ (3.5)
Case 2: Let $\beta_{1}=j\in\{0,1,\cdots,m-2\}$ be fixed.
  • (2a)
    There is a unique $\beta_{2}(j)\in{\mathbb{Z}}$ so that $\kappa(j,\beta_{2}(j))\in\{0,1,\cdots,m-2\}$.
    Explicitly,
    $\beta_{2}(j)=2n-L(j),\qquad\kappa(j,\beta_{2}(j))=m+n-1+nj-mL(j),$ (3.6)
    so each $\beta_{2}(j)\in\{n,n+1,\cdots,2n-1\}$. Distinct choices of $j_{1},j_{2}\in\{0,1,\cdots,m-2\}$ yield $\kappa(j_{1},\beta_{2}(j_{1}))\neq\kappa(j_{2},\beta_{2}(j_{2}))$, meaning that $\kappa(j,\beta_{2}(j))$ assumes each value in $\{0,1,\cdots,m-2\}$ exactly once as $j$ ranges in $\{0,1,\cdots,m-2\}$.
  • (2b)
    There is a unique $\tilde{\beta}_{2}(j)\in{\mathbb{Z}}$ so that $\kappa(j,\tilde{\beta}_{2}(j))\in\{m,m+1,\cdots,2m-2\}$.
    Explicitly,
    $\tilde{\beta}_{2}(j)$ $=2n+1-L(j),\qquad\kappa(j,\tilde{\beta}_{2}(j))$ $=2m+n-1+nj-mL(j),$ (3.7)
    $=\beta_{2}(j)+1,\qquad$ $=\kappa(j,\beta_{2}(j))+m.$
    so each $\tilde{\beta}_{2}(j)\in\{n+1,n+2,\cdots,2n\}$. Distinct choices of $j_{1},j_{2}\in\{0,1,\cdots,m-2\}$ yield $\kappa(j_{1},\tilde{\beta}_{2}(j_{1}))\neq\kappa(j_{2},\tilde{\beta}_{2}(j_{2}))$, meaning that $\kappa(j,\tilde{\beta}_{2}(j))$ assumes each value in $\{m,m+1,\cdots,2m-2\}$ exactly once as $j$ ranges in $\{0,1,\cdots,m-2\}$.
Case 3: Let $\beta_{1}=j\in\{m,m+1,\cdots,2m-2\}$ be fixed.
  • (3a)
    There is a unique $\beta_{2}(j)\in{\mathbb{Z}}$ so that $\kappa(j,\beta_{2}(j))\in\{0,1,\cdots,m-2\}$.
    Explicitly,
    $\beta_{2}(j)=2n-L(j),\qquad\kappa(j,\beta_{2}(j))=m+n-1+nj-mL(j),$ (3.8)
    so each $\beta_{2}(j)\in\{0,1,\cdots,n-1\}$. Distinct choices of $j_{1},j_{2}\in\{m,m+1,\cdots,2m-2\}$ yield $\kappa(j_{1},\beta_{2}(j_{1}))\neq\kappa(j_{2},\beta_{2}(j_{2}))$, meaning that $\kappa(j,\beta_{2}(j))$ assumes each value in $\{0,1,\cdots,m-2\}$ exactly once as $j$ ranges in $\{m,m+1,\cdots,2m-2\}$.
  • (3b)
    There is a unique $\tilde{\beta}_{2}(j)\in{\mathbb{Z}}$ so that $\kappa(j,\tilde{\beta}_{2}(j))\in\{m,m+1,\cdots,2m-2\}$.
    Explicitly,
    $\tilde{\beta}_{2}(j)$ $=2n+1-L(j),\qquad\kappa(j,\tilde{\beta}_{2}(j))$ $=2m+n-1+nj-mL(j),$ (3.9)
    $=\beta_{2}(j)+1,\qquad$ $=\kappa(j,\beta_{2}(j))+m.$
    so each $\tilde{\beta}_{2}(j)\in\{1,2,\cdots,n\}$. Distinct choices of $j_{1},j_{2}\in\{m,m+1,\cdots,2m-2\}$ yield $\kappa(j_{1},\tilde{\beta}_{2}(j_{1}))\neq\kappa(j_{2},\tilde{\beta}_{2}(j_{2}))$, meaning that $\kappa(j,\tilde{\beta}_{2}(j))$ assumes each value in $\{m,m+1,\cdots,2m-2\}$ exactly once as $j$ ranges in $\{m,m+1,\cdots,2m-2\}$.
Proof.
Statement $(0)$ is clear from the definition of $D_{m}(\beta)$ given in (1.10), but it sets the stage for the three subsequent cases.
Case $1$ concerns the situation when either $\beta_{1}$ or $\kappa(\beta_{1},\beta_{2})$ is equal to $m-1$, that is, the input value where the function $D_{m}$ peaks.
For $(1a)$, let $\beta_{1}=m-1$. If $\beta_{2}\in{\mathbb{Z}}$ is chosen so that $20\leq\kappa(m-1,\beta_{2})\leq 2m-$ 2, definition (3.2) shows
$n-1+\frac{1}{m}\leq\beta_{2}\leq n+1-\frac{1}{m}.$
This forces $\beta_{2}=n$, and a computation shows $\kappa(m-1,n)=m-1$.
For $(1b)$, assume $\beta_{1},\beta_{2}$ are chosen so that $\kappa(\beta_{1},\beta_{2})=m-1$. This means
$m(\beta_{2}-2n)=-n(\beta_{1}+1)\qquad\Longrightarrow\qquad\beta_{1}\equiv-1\mod m,$
since $\gcd(m,n)=1$. Since $20\leq\beta_{1}\leq 2m-$ 2, this forces $\beta_{1}=m-1$, which implies that $\beta_{2}=n$. This concludes case $1$.
In Case 2, we let $\beta_{1}=j\in\{0,1,\cdots,m-2\}$, i.e., the inputs where $D_{m}$ is ascending, up to (but not including) the peak at $j=m-1$.
For $(2a)$, assume that $\beta_{2}$ is chosen so that $20\leq\kappa(j,\beta_{2})\leq m-$ 2. This means
$2mn+1-m-n-nj\leq m\beta_{2}\leq 2mn-1-n-nj$ (3.10)
In the case $m=2$ (which forces $n=1$), we immediately have $j=0$, $\beta_{2}=1$ and $\kappa(0,1)=0$. For integers $m\geq 3$, (3.10) is a closed interval of length $m-2$ with integer endpoints. We claim this interval must contain a multiple of $m$. If not, the left endpoint would satisfy
$2mn+1-m-n-nj\equiv 1\mod m.$
But this would then necessitate
$-n(j+1)\equiv 0\mod m\qquad\Longrightarrow\qquad j\equiv-1\mod m,$
which is impossible since $j\in\{0,1\cdots,m-2\}$. Thus, for each $j$ in this range, there is a unique $\beta_{2}=\beta_{2}(j)\in{\mathbb{Z}}$ satisfying (3.10), given by
$\beta_{2}(j)=\left\lfloor 2n-\frac{1+n(j+1)}{m}\right\rfloor=2n-L(j),$ (3.11)
where $L(j)$ is defined by (3.5). Since $\frac{n}{m}<1$, $L(j)$ is a non-decreasing function of $j$. Inspection shows that $L(0)=1$ and $L(m-2)=n$, with each integer value between $1$ and $n$ attained at least once. This means that $\beta_{2}(0)=2n-1$ and that $\beta_{2}(m-2)=n$, with every integer value between attained by some $\beta_{2}(j)$ at least once.
Direct evaluation from (3.2) shows
$\kappa(j,\beta_{2}(j))=m+n-1+nj-mL(j).$
Now if $j_{1},j_{2}\in\{0,1,\cdots,m-2\}$ are chosen so that $\kappa(j_{1},\beta_{2}(j_{1}))=\kappa(j_{2},\beta_{2}(j_{2}))$, we must have that
$n(j_{1}-j_{2})=m(L(j_{1})-L(j_{2}))\qquad\Longrightarrow\qquad j_{1}\equiv j_{2}\mod m,$
which is only possible if $j_{1}=j_{2}$. We thus conclude that $\kappa(j,\beta_{2}(j))$ attains each value in $\{0,1,\cdots,m-2\}$ exactly once as $j$ varies throughout the same range.
Part $(2b)$ is easily deduced from $(2a)$. Take $\tilde{\beta}_{2}(j)=\beta_{2}(j)+1$, where $\beta_{2}(j)$ is given by (3.11), and see from (3.2) that
$\kappa(j,\tilde{\beta}_{2}(j))$ $=m\tilde{\beta}_{2}(j)+njm+n-1-2mn$
$=m+m\beta_{2}(j)+njm+n-1-2mn$
$=m+\kappa(j,\beta_{2}(j)).$
Now since $\{m,m+1,\cdots,2m-2\}$ is just a shift of $\{0,1,\cdots,m-2\}$ by $m$ units, all claims made in $(2b)$ are equivalent to those in $(2a)$. This concludes case 2.
Case 3 is handled similarly, but now $\beta_{1}=j$ ranges in $\{m,m+1,\cdots,2m-2\}$, i.e., the range of values (not including the peak at $j=m-1$) on which $D_{m}$ is descending.
For $(3a)$, assume $\beta_{2}$ is chosen so that $20\leq\kappa(j,\beta_{2})\leq m-$ 2. This yields the same inequalities that appeared in (3.10), only now with a different range of $j$ values. When $m=2$ (which forces $n=1$), we immediately have that $j=2$, so $\beta_{2}=0$ and $\kappa(2,0)=0$. For integers $m\geq 3$, (3.10) is a closed interval of length $m-2$ with integer endpoints. Following reasoning identical to that given above for $(2a)$, we see that for each $j\in\{m,m+1,\cdots,2m-2\}$, there is a unique $\beta_{2}=\beta_{2}(j)\in{\mathbb{Z}}$ satisfying (3.10). Once again $\beta_{2}(j)=2n-L(j)$, the same formula provided in (3.11) for $j\in\{0,1,\cdots,m-2\}$. Inspection shows that $\beta_{2}(m)=n-1$, and that $\beta_{2}(2m-2)=0$, and reasoning identical to that given above shows that $\beta_{2}(j)$ attains each integer between $n-1$ and 00 at least once.
Direct evaluation from (3.2) again shows that $\kappa(j,\beta_{2}(j))=m+n-1+nj-mL(j)$, and identical reasoning to that given above shows that if $\kappa(j_{1},\beta_{2}(j_{1}))=\kappa(j_{2},\beta_{2}(j_{2}))$ for $j_{1},j_{2}\in\{m,m+1,\cdots,2m-2\}$, then $j_{1}=j_{2}$. We thus conclude that $\kappa(j,\beta_{2}(j))$ attains each value in $\{0,1,\cdots,m-2\}$ exactly once as $j$ varies throughout $\{m,m+1,\cdots,2m-2\}$.
Part $(3b)$ is easily deduced from $(3a)$ in a way that parallels how $(2b)$ was deduced from $(2a)$. We once again set $\tilde{\beta}_{2}(j)=\beta_{2}(j)+1$, which is seen to imply that $\kappa(j,\tilde{\beta}_{2}(j))=m+\kappa(j,\beta_{2}(j))$. As above, all claims made in $(3b)$ equivalent to those made in $(3a)$. This concludes case 3 and completes the proof. ∎
3.2. Explicit determination of $P_{m,n}(s,t)$
We produce an explicit formula for the numerator of the Bergman kernel in terms of expressions appearing in Proposition 3.4; the quantities $\kappa(j,\beta_{2}(j))$ and $\kappa(j,\tilde{\beta}_{2}(j))$ will be abbreviated as follows:
$\kappa(j)$ $:=\kappa(j,\beta_{2}(j))=m+n-1+nj-mL(j),$ (3.12a)
$\tilde{\kappa}(j)$ $:=\kappa(j,\tilde{\beta}_{2}(j))=2m+n-1+nj-mL(j)=m+\kappa(j).$ (3.12b)
We first record some properties of $L(j)$ and $\kappa(j)$ that will be crucial later.
Lemma 3.13.
The functions $L$ and $\kappa$, defined in (3.5) and (3.12a), respectively, satisfy the following
  1. (1)
    $L(j+m)=L(j)+n$.
  2. (2)
    $\kappa(j+m)=\kappa(j)$.
  3. (3)
    For $j\in\{0,1,\cdots,m-2\}$, $L(j)+L(m-2-j)=n+1$.
  4. (4)
    For $j\in\{0,1,\cdots,m-2\}$, $\kappa(j)+\kappa(m-2-j)=m-2$.
Proof.
(1) and (2) follow from trivial computation. For (3), observe that (1) implies $L(j)+L(m-2-j)=n+L(j)+L(-2-j)$.
By definition, $L(j)+L(-2-j)=$
$=\left\lceil\frac{1+n(j+1)}{m}\right\rceil+\left\lceil\frac{1-n(j+1)}{m}\right\rceil=\left\lceil\frac{n(j+1)+1}{m}\right\rceil-\left\lfloor\frac{n(j+1)-1}{m}\right\rfloor.$ (3.14)
For $j\in\{0,1,\cdots,m-2\}$, write $n(j+1)=q(j)m+r(j)$, where $q(j)\in{\mathbb{Z}}$ is the quotient and $r(j)$ is the remainder. We claim $11\leq r(j)\leq m-$ 1 for this range of $j$. Indeed, since $\gcd(m,n)=1$, we have that $r(j)$ assumes each value in the set $\{1,2,\cdots,m-1\}$ exactly once. Thus,
$\frac{n(j+1)+1}{m}=q(j)+\frac{r(j)+1}{m},\qquad\frac{n(j+1)-1}{m}=q(j)+\frac{r(j)-1}{m},$
and so
$\left\lceil\frac{n(j+1)+1}{m}\right\rceil=q(j)+1,\qquad\left\lfloor\frac{n(j+1)-1}{m}\right\rfloor=q(j).$
Now by (3.14) we see that for $j\in\{0,1,\cdots,m-2\}$, $L(-2-j)+L(j)=1$, proving (3).
For (4), notice that
$\kappa(m-2-j)+\kappa(j)$ $=2m+2n-2+n(m-2-j)-mL(m-2-j)+nj-mL(j)$
$=2m-2+nm-m(L(m-2-j)+L(j))$
$=2m-2+nm-m(n+1)=m-2,$ (3.15)
completing the proof. ∎
We are ready to establish the Bergman kernel formula.
Proof of Theorem 1.3.
We use parts of Proposition 3.4 to decompose the polynomial
$P_{m,n}(s,t)=\sum_{\beta\in{\mathbb{Z}}^{2}}D_{m}(\beta_{1})D_{m}(m\beta_{2}+n\beta_{2}+m+n-1-2mn)s^{\beta_{1}}t^{\beta_{2}}.$
  • Parts (1a) and (1b) together show that the two possible options in case 1 ($\beta_{1}=m-1$ or $\kappa(\beta_{1},\beta_{2})=m-1$) in fact both correspond to the same, uniquely determined $\beta$-value: $(\beta_{1},\beta_{2})=(m-1,n)$. This corresponds to a single monomial term in the sum above, which we now name. Recalling (1.10), define
    $p_{0}(s,t)$ $:=D_{m}(\beta_{1})D_{m}(\kappa(\beta_{1},\beta_{2}))s^{\beta_{1}}t^{\beta_{2}}\Big{|}_{(\beta_{1},\beta_{2})=(m-1,n)}$
    $=D_{m}(m-1)D_{m}(m-1)s^{m-1}t^{n}$
    $=m^{2}s^{m-1}t^{n}.$
  • Part (2a) considers $\beta_{1}=j\in\{0,1,\cdots,m-2\}$ and shows the integer $\beta_{2}(j)=2n-L(j)$ is uniquely determined so that $\kappa(j):=\kappa(j,\beta_{2}(j))\in\{0,1,\cdots,m-2\}$. Recalling (1.10), define the polynomial corresponding to these indices:
    $p_{1}(s,t)$ $:=\sum_{j=0}^{m-2}D_{m}(j)D_{m}(\kappa(j))s^{j}t^{\beta_{2}(j)}=\sum_{j=0}^{m-2}(j+1)(\kappa(j)+1)s^{j}t^{2n-L(j)}.$
  • Part (2b) considers $\beta_{1}=j\in\{0,1,\cdots,m-2\}$ and shows the integer $\tilde{\beta}_{2}(j)=2n+1-L(j)$ is uniquely determined so that $\tilde{\kappa}(j):=\kappa(j,\tilde{\beta}_{2}(j))=m+\kappa(j)\in\{m,m+1,\cdots,2m-2\}$. Recalling (1.10), we define the polynomial corresponding to these indices:
    $p_{2}(s,t)$ $:=\sum_{j=0}^{m-2}D_{m}(j)D_{m}(m+\kappa(j))s^{j}t^{\tilde{\beta}_{2}(j)}$
    $=\sum_{j=0}^{m-2}(j+1)(m-\kappa(j)-1)s^{j}t^{2n+1-L(j)}.$
  • Part (3a) considers $\beta_{1}=j\in\{m,m+1,\cdots,2m-2\}$ and shows the integer $\beta_{2}(j)=2n-L(j)$ is uniquely determined so that $\kappa(j):=\kappa(j,\beta_{2}(j))\in\{0,1,\cdots,m-2\}$. Recalling (1.10), define the polynomial corresponding to these indices:
    $p_{3}(s,t)$ $:=\sum_{j=m}^{2m-2}D_{m}(j)D_{m}(\kappa(j))s^{j}t^{\beta_{2}(j)}$
    $=\sum_{j=m}^{2m-2}(2m-j-1)(\kappa(j)+1)s^{j}t^{2n-L(j)}=\sum_{j=0}^{m-2}(m-j-1)(\kappa(j)+1)s^{j+m}t^{n-L(j)}.$
    In the last step, we re-indexed the sum using Lemma 3.13 to see that $\kappa(j+m)=\kappa(j)$ and $L(j+m)=L(j)-n$.
  • Part (3b) considers $\beta_{1}=j\in\{m,m+1,\cdots,2m-2\}$ and shows the integer $\tilde{\beta}_{2}(j)=2n+1-L(j)$ is uniquely determined so that $\tilde{\kappa}(j):=\kappa(j,\tilde{\beta}_{2}(j))=m+\kappa(j)\in\{m,m+1,\cdots,2m-2\}$. Recalling (1.10), define the polynomial corresponding to these indices:
    $p_{4}(s,t)$ $:=\sum_{j=m}^{2m-2}D_{m}(j)D_{m}(m+\kappa(j))s^{j}t^{\tilde{\beta}_{2}(j)}$
    $=\sum_{j=m}^{2m-2}(2m-j-1)(m-\kappa(j)-1)s^{j}t^{2n+1-L(j)}$
    $=\sum_{j=0}^{m-2}(m-j-1)(m-\kappa(j)-1)s^{j+m}t^{n+1-L(j)}.$
As above, we re-indexed the last sum using Lemma 3.13. By Proposition 3.4 we have exhausted all indices $(\beta_{1},\beta_{2})$ corresponding to non-zero coefficients in the sum defining $P_{m,n}(s,t)$ meaning the proof is now complete. ∎
3.3. Palindromic polynomials
In this section we restrict the numerator polynomial of the Bergman kernel to the line $\{s=t\}$. As the function $s\mapsto P_{m,n}(s,s)$ has a root of order $12n-$ 1 at $s=0$ (this is shown below), we are led to define a new polynomial
$Q_{m,n}(s):=s^{1-2n}P_{m,n}(s,s).$ (3.16)
Recalling the discussion in Section 1.2.1, we observe that $Q_{m,n}$ has roots in the punctured unit disc ${\mathbb{D}}^{*}\subset{\mathbb{C}}$ if and only if $P_{m,n}$ has roots in the punctured disc $\Delta^{*}=\{(s,s):s\in{\mathbb{D}}^{*}\}\subset{\mathbb{H}}_{m/n}$.
Our main result this section is that $Q_{m,n}$ is a palindromic polynomial of degree $2m-2n$. Recall that a degree $2k$ polynomial
$f(s)=\sum_{j=0}^{2k}\alpha_{j}s^{j}$
is called palindromic when its coefficients satisfy $\alpha_{j}=\alpha_{2k-j}$ for $0\leq j\leq k$.
With given $f$ as above, it is palindromic if and only if it satisfies
$f(s)=s^{2k}f(s^{-1}).$ (3.17)
Equation (3.17) shows that $\0s\in{\mathbb{C}}\backslash\{0\}$ is a root of the palindromic polynomial $f$ if and only if $s^{-1}$ is also a root. This immediately implies the following:
Proposition 3.18.
Suppose $f$ is a palindromic polynomial of degree $2k$. Then either
  1. (1)
    $f$ has all its roots on the unit circle, or
  2. (2)
    $f$ has an even number of roots (counting multiplicities) off the unit circle, exactly half of which lie inside ${\mathbb{D}}^{*}$.
In particular, if $f$ is non-vanishing on the unit circle, then it has exactly $k$ roots inside ${\mathbb{D}}^{*}$.
Following (1.4b), we express $Q_{m,n}$ as the sum
$Q_{m,n}(s)=q_{0}(s)+q_{1}(s)+q_{2}(s)+q_{3}(s)+q_{4}(s),$ (3.19)
where each
$q_{j}(s):=s^{1-2n}p_{j}(s,s).$
The formulas for the $p_{j}$ given in Theorem 1.3 then yield
$q_{0}(s)=m^{2}s^{m-n},$ (3.20a)
$q_{1}(s)=\sum_{j=0}^{m-2}(j+1)(\kappa(j)+1)s^{j-L(j)+1},$ (3.20b)
$q_{2}(s)=\sum_{j=0}^{m-2}(j+1)(m-\kappa(j)-1)s^{j-L(j)+2},$ (3.20c)
$q_{3}(s)=\sum_{j=0}^{m-2}(m-j-1)(\kappa(j)+1)s^{j-L(j)+m-n+1},$ (3.20d)
$q_{4}(s)=\sum_{j=0}^{m-2}(m-j-1)(m-\kappa(j)-1)s^{j-L(j)+m-n+2}.$ (3.20e)
The $q_{j}$ satisfy interesting symmetries:
Proposition 3.21.
The $q_{j}$ are polynomials with positive coefficients, satisfying
$s^{2m-2n}q_{0}(s^{-1})=q_{0}(s),$ (3.22a)
$s^{2m-2n}q_{1}(s^{-1})=q_{4}(s),$ (3.22b)
$s^{2m-2n}q_{2}(s^{-1})=q_{3}(s).$ (3.22c)
Consequently, $Q_{m,n}$ is a palindromic polynomial of degree $2m-2n$ with positive coefficients.
Proof.
Each coefficient inside the sums (3.20b)-(3.20e) is seen to be positive by recalling Proposition 3.4 parts $(2a)$ and $(3a)$, that $\kappa(j)$ assumes each value in $\{0,1,\cdots,m-2\}$ exactly once as $j$ ranges in $\{0,1,\cdots,m-2\}$. (Recall that in (3.12a) we simplified notation by writing $\kappa(j)$ for what was originally defined by $\kappa(j,\beta_{2}(j))$).
Next we show that each $q_{j}$ is a polynomial, i.e., that only non-negative powers of $s$ appear. Clearly $q_{0}$ is a polynomial since we assume $m>n$. Now notice that in the sums defining $q_{1},q_{2},q_{3},q_{4}$ above, the exponent on $q_{1}$ in (3.20b) is smaller than the other exponents. We write this exponent as
$E(j):=j-L(j)+1.$ (3.23)
We now show this is non-negative for all $j=0,1,\cdots,m-2$.
First, we claim $E(j)$ is a non-decreasing function on the integers, i.e., $E(j+1)-E(j)\geq 0$. To see this, observe
$E(j+1)-E(j)=1-\left(\left\lceil\frac{1+n(j+2)}{m}\right\rceil-\left\lceil\frac{1+n(j+1)}{m}\right\rceil\right)$
Looking at the terms inside the ceiling brackets, we have
$\frac{1+n(j+2)}{m}-\frac{1+n(j+1)}{m}=\frac{n}{m}<1,$
meaning that for any $j$, the difference in parentheses above takes one of two possible values:
$\left\lceil\frac{1+n(j+2)}{m}\right\rceil-\left\lceil\frac{1+n(j+1)}{m}\right\rceil\in\{0,1\}.$
This means that for each $j$, $E(j+1)-E(j)\geq 0$. It is thus sufficient to check the value at $j=0$, the smallest $j$-value appearing in the sum:
$E(0)=1-L(0)=1-\left\lceil\frac{1+n}{m}\right\rceil=0,$ (3.24)
since $0<1+n\leq m$. This shows that each $q_{1}$ is a polynomial with a non-zero constant term; consequently $q_{2},q_{3},q_{4}$ are also polynomials.
Now we verify the identities in (3.22). First,
$s^{2m-2n}q_{0}(s^{-1})=s^{2m-2n}m^{2}s^{n-m}=m^{2}s^{m-n}=q_{0}(s),$
confirming (3.22a). Next consider
$s^{2m-2n}q_{1}(s^{-1})=\sum_{j=0}^{m-2}(j+1)(\kappa(j)+1)s^{2m-2n+L(j)-j-1},$ (3.25)
and re-index the sum by setting $j=m-2-i$. By Lemma 3.13, we have that $\kappa(m-2-i)=m-2-\kappa(i)$ and $L(m-2-i)=n+1-L(i)$, and so
[…]
$=\sum_{i=0}^{m-2}(m-1-i)(m-1-\kappa(i))s^{i-L(i)+m-n+2}=q_{4}(s),$
confirming (3.22b). Following the same lines of reasoning, now consider
$s^{2m-2n}q_{2}(s^{-1})$ $=\sum_{j=0}^{m-2}(j+1)(m-\kappa(j)-1)s^{2m-2n+L(j)-j-2}$
$=\sum_{i=0}^{m-2}(m-1-i)(m-\kappa(m-2-i)-1)s^{2m-2n+L(m-2-i)-m+2+i-2}$
$=\sum_{i=0}^{m-2}(m-1-i)(\kappa(i)+1)s^{i-L(i)+m-n+1}=q_{3}(s),$
confirming (3.22c). Finally, we show $Q_{m,n}(s)$ is palindromic of degree $2m-2n$:
$s^{2m-2n}Q_{m,n}(s^{-1})$ $=s^{2m-2n}\big{(}q_{0}(s^{-1})+q_{1}(s^{-1})+q_{2}(s^{-1})+q_{3}(s^{-1})+q_{4}(s^{-1})\big{)}$
$=q_{0}(s)+q_{4}(s)+q_{3}(s)+q_{2}(s)+q_{1}(s)$
$=Q_{m,n}(s),$
as (3.22b) and (3.22c) also imply $s^{2m-2n}q_{4}(s^{-1})=q_{1}(s)$ and $s^{2m-2n}q_{3}(s^{-1})=q_{2}(s)$. ∎
4. The Lu Qi-Keng problem
We use the work of the previous section to prove the existence of zeros of Bergman kernels on ${\mathbb{H}}_{\gamma}$ for infinitely many rationals $\gamma=\frac{m}{n}\in(1,2)$.
4.1. Case $m-n=1$
$\big{(}\frac{m}{n}=\frac{2}{1},\frac{3}{2},\frac{4}{3},\frac{5}{4},\frac{6}{5},\frac{7}{6}\cdots\big{)}$.
Proposition 4.1.
Let us write $(m,n)=(\ell+1,\ell)$, $\ell\in{\mathbb{Z}}^{+}$. The polynomial $Q_{\ell+1,\ell}$ takes the form
$Q_{\ell+1,\ell}(s)$ $=\alpha_{0}(\ell)s^{2}+\alpha_{1}(\ell)s+\alpha_{0}(\ell),$
where
$\alpha_{0}(\ell)$ $=\tfrac{1}{6}\ell(\ell+1)(\ell+2),$
$\alpha_{1}(\ell)$ $=\tfrac{1}{3}(\ell+1)(2\ell^{2}+4\ell+3).$
Proof.
For $(m,n)=(\ell+1,\ell)$ let us first calculate the quantities $L(j)$ and $\kappa(j)$ from (1.4h). For the range of $j$ appearing in the sums defining $q_{1},q_{2},q_{3},q_{4}$ $(0\leq j\leq m-2=\ell-1)$,
$L(j)=\left\lceil\frac{1+\ell(j+1)}{\ell+1}\right\rceil=j+1.$ (4.2)
A calculation then shows $\kappa(j)=\ell-j-1$ and thus equations (3.20a)–(3.20e) give
$q_{0}(s)=(\ell+1)^{2}s,\qquad q_{1}(s)=\sum_{j=0}^{\ell-1}(j+1)(\ell-j),\qquad q_{2}(s)=s\sum_{j=0}^{\ell-1}(j+1)^{2}$
$q_{3}(s)=s\sum_{j=0}^{\ell-1}(\ell-j)^{2},\qquad q_{4}(s)=s^{2}\sum_{j=0}^{\ell-1}(\ell-j)(j+1).$
Now add these polynomials:
$Q_{\ell+1,\ell}(s,s)$ $=q_{0}(s)+q_{1}(s)+q_{2}(s)+q_{3}(s)+q_{4}(s)$
$=(1+s^{2})\left(\sum_{j=0}^{\ell-1}(j+1)(\ell-j)\right)+s\left(2\sum_{j=0}^{\ell-1}(j+1)^{2}+(\ell+1)^{2}\right)$
$=(1+s^{2})\alpha_{0}(\ell)+s\alpha_{1}(\ell),$
where the formulas for $\alpha_{0}(\ell)$ and $\alpha_{1}(\ell)$ as cubics in $\ell$ follow from well known sums of powers formulas (see, e.g., [18]). ∎
Proposition 4.3.
$Q_{\ell+1,\ell}(s)$ has a unique root inside the punctured unit disc ${\mathbb{D}}^{*}$.
Proof.
Let us factor
$Q_{\ell+1,\ell}(s)$ $=\alpha_{0}(\ell)s^{2}+\alpha_{1}(\ell)s+\alpha_{0}(\ell)$
$=\alpha_{0}(\ell)\big{(}s-r_{1}(\ell)\big{)}\big{(}s-r_{2}(\ell)\big{)},$ (4.4)
where $r_{1}(\ell)$ and $r_{2}(\ell)$ denote the two roots. Observe that
$\alpha_{1}(\ell)=\tfrac{1}{3}(\ell+1)(2\ell^{2}+4\ell+3)>\tfrac{1}{3}(\ell+1)(\ell^{2}+2\ell)=2\alpha_{0}(\ell),$ (4.5)
which shows the roots $r_{1}(\ell)$ and $r_{2}(\ell)$ are real-valued and distinct. Also observe from (4.4) that $r_{1}(\ell)\cdot r_{2}(\ell)=1$, which implies the roots have the same sign. Thus, one of them (say $r_{1}(s)$) lies inside ${\mathbb{D}}^{*}$ while $r_{2}(\ell)$ lies outside $\overline{{\mathbb{D}}}$. (It is clear that the roots must be negative, as the coefficient functions $\alpha_{0}(\ell)$ and $\alpha_{1}(\ell)$ are always positive.) ∎
Remark 4.6.
We can find the above roots explicitly. The quadratic formula shows
$r_{1}(\ell)=-\left(\frac{2\ell^{2}+4\ell+3}{\ell^{2}+2\ell}\right)+\sqrt{\left(\frac{2\ell^{2}+4\ell+3}{\ell^{2}+2\ell}\right)^{2}-1}.$
Now observe that as $\ell\to\infty$, this root remains inside the punctured disc:
$\lim_{\ell\to\infty}r_{1}(\ell)=-2+\sqrt{3}\in{\mathbb{D}}^{*}.$
Corollary 4.7.
The Bergman kernel of ${\mathbb{H}}_{\frac{\ell+1}{\ell}}$ has zeros inside ${\mathbb{H}}_{\frac{\ell+1}{\ell}}\times{\mathbb{H}}_{\frac{\ell+1}{\ell}}$ for every $\ell\in{\mathbb{Z}}^{+}$.
Proof.
When $\Omega$ is Reinhardt, Corollary 2.12 shows the Bergman kernel $K_{\Omega}(z,w)$ has zeros inside $\Omega\times\Omega$ if and only if the corresponding holomorphic factor $k_{\Omega}(s,t)$ has zeros on the domain $\Omega^{(2)}$. But the domains satisfy ${\mathbb{H}}_{\gamma}^{(2)}={\mathbb{H}}_{\gamma}$ (Proposition 2.13), and (1.4a) gives
$k_{\frac{\ell+1}{\ell}}(s,t)=\frac{1}{\pi^{2}(\ell+1)}\cdot\frac{P_{\ell+1,\ell}(s,t)}{(1-s^{2})^{2}(t^{\ell}-s^{\ell+1})^{2}},$
where the denominator is non-vanishing for $(s,t)\in{\mathbb{H}}_{\frac{\ell+1}{\ell}}$. Proposition 4.13 shows that $Q_{\ell+1,\ell}$ has a root in ${\mathbb{D}}^{*}$, and (3.16) shows that this root corresponds to a root of $P_{\ell+1,\ell}$ restricted to the punctured disc $\Delta^{*}=\{(s,s)\in{\mathbb{C}}^{2}:s\in{\mathbb{D}}^{*}\}\subset{\mathbb{H}}_{\frac{\ell+1}{\ell}}$. ∎
4.2. Case $m-n=2$
$\big{(}\frac{m}{n}=\frac{3}{1},\frac{5}{3},\frac{7}{5},\frac{9}{5},\frac{11}{9},\frac{13}{11},\cdots\big{)}$.
Proposition 4.8.
Let $(m,n)=(2\ell+1,2\ell-1)$, $\ell\in{\mathbb{Z}}^{+}$. The polynomial $Q_{2\ell+1,2\ell-1}$ takes the form
$Q_{2\ell+1,2\ell-1}(s)$ $=\alpha_{0}(\ell)(1+s^{4})+\alpha_{1}(\ell)(s+s^{3})+\alpha_{2}(\ell)s^{2},$
where
$\alpha_{0}(\ell)$ $=\tfrac{1}{6}\ell(\ell+1)(2\ell+1),$
$\alpha_{1}(\ell)$ $=\ell(\ell+1)(2\ell+1),$
$\alpha_{2}(\ell)$ $=\tfrac{1}{3}(2\ell+1)(5\ell^{2}+5\ell+3).$
Proof.
Again we compute $L(j)$ and $\kappa(j)$ from (1.4h), and now the relevant $j$ values satisfy $10\leq j\leq m-2=2\ell-$ 1.
$L(j)=\left\lceil\frac{1+(2\ell-1)(j+1)}{2\ell+1}\right\rceil$ $=j+1-\left\lfloor\frac{2(j+1)-1}{2\ell+1}\right\rfloor$
=$=\begin{cases}j+1,&\quad 0\leq j\leq\ell-1,\\ j,&\quad\ell\leq j\leq 2\ell-1.\end{cases}$ (4.9)
A short calculation shows
$\kappa(j)=\begin{cases}2\ell-2-2j,&\quad 0\leq j\leq\ell-1,\\ 4\ell-1-2j,&\quad\ell\leq j\leq 2\ell-1.\end{cases}$ (4.10)
Equations (3.20a)–(3.20c) now give
$q_{0}(s)$ $=(2\ell+1)^{2}s^{2},$ (4.11)
$q_{1}(s)$ $=\sum_{j=0}^{2\ell-1}(j+1)(\kappa(j)+1)s^{j-L(j)+1}$
$=\sum_{j=0}^{\ell-1}(j+1)(2\ell-1-2j)s^{j-(j+1)+1}+\sum_{j=\ell}^{2\ell-1}(j+1)(4\ell-2j)s^{j-j+1}$
$=\left(\frac{\ell(\ell+1)(2\ell+1)}{6}\right)+\left(\frac{2\ell(\ell+1)(2\ell+1)}{3}\right)s,$ (4.12)
$q_{2}(s)$ $=\sum_{j=0}^{2\ell-1}(j+1)(2\ell-\kappa(j))s^{j-L(j)+2}$
$=2s\sum_{j=0}^{\ell-1}(j+1)^{2}+s^{2}\sum_{j=\ell}^{2\ell-1}(j+1)(2j+1-2\ell)$
$=\left(\frac{\ell(\ell+1)(2\ell+1)}{3}\right)s+\left(\frac{\ell(2\ell+1)(5\ell-1)}{6}\right)s^{2}.$
By Proposition 3.21 we see
$q_{3}(s)$ $=\left(\frac{\ell(\ell+1)(2\ell+1)}{3}\right)s^{3}+\left(\frac{\ell(2\ell+1)(5\ell-1)}{6}\right)s^{2},$
$q_{4}(s)$ $=\left(\frac{\ell(\ell+1)(2\ell+1)}{6}\right)s^{4}+\left(\frac{2\ell(\ell+1)(2\ell+1)}{3}\right)s^{3}.$
Now add $q_{0}+q_{1}+q_{2}+q_{3}+q_{4}$ to confirm $Q_{2\ell+1,2\ell-1}(s)$ takes the form given in the statement of the theorem. ∎
Proposition 4.13.
$Q_{2\ell+1,2\ell-1}(s)$ has exactly two roots inside the punctured unit disc ${\mathbb{D}}^{*}$.
Proof.
To locate the roots of $Q_{2\ell+1,2\ell-1}(s)$, we start with the formula in Proposition 4.8 and factor out the coefficient of $(1+s^{4})$:
$Q_{2\ell+1,2\ell-1}(s)$ $=\frac{\ell(\ell+1)(2\ell+1)}{6}\left(1+6s+\left(10+\frac{6}{\ell^{2}+\ell}\right)s^{2}+6s^{3}+s^{4}\right)$
$:=\frac{\ell(\ell+1)(2\ell+1)}{6}s^{2}R_{\ell}(s),$ (4.14)
where
$R_{\ell}(s)=s^{-2}+6s^{-1}+\left(10+\frac{6}{\ell^{2}+\ell}\right)+6s+s^{2}.$ (4.15)
The roots of $Q_{2\ell+1,2\ell-1}$ are exactly the roots of $R_{\ell}(s)$. We now show that $R_{\ell}(s)$ is non-vanishing on the unit circle.
Suppose the contrary, that $s=e^{i\theta}$ is a root for some $\theta\in{\mathbb{R}}$. Then
$R_{\ell}(e^{i\theta})=4\cos^{2}\theta+12\cos\theta+\left(8+\frac{6}{\ell^{2}+\ell}\right)=0,$ (4.16)
which is equivalent to saying there exists $x=\cos\theta\in[-1,1]$ such that
$G_{\ell}(x):=4x^{2}+12x+\beta(\ell)=0,\qquad\mathrm{where}\qquad\beta(\ell)=\left(8+\frac{6}{\ell^{2}+\ell}\right).$ (4.17)
But the roots of $G_{\ell}(s)$ can be calculated explicitly:
$r_{1}(\ell)=-\frac{3}{2}+\frac{1}{2}\sqrt{9-\beta(\ell)},\qquad r_{2}(\ell)=-\frac{3}{2}-\frac{1}{2}\sqrt{9-\beta(\ell)}.$ (4.18)
Note that $\beta$ is a decreasing function with $\beta(1)=11$, $\beta(2)=9$ and $\lim_{\ell\to\infty}\beta(\ell)=8$.
When $\ell=1$, the roots are the complex numbers
$r_{1}(1)=-\frac{3}{2}+i\frac{\sqrt{2}}{2},\qquad r_{2}(1)=-\frac{3}{2}-i\frac{\sqrt{2}}{2},$
which are clearly not in the interval $[-1,1]$. For $\ell\geq 2$, the roots are real numbers given by
$r_{1}(\ell)\in\left[-\frac{3}{2},-1\right),\qquad r_{2}(\ell)\in\left(-2,-\frac{3}{2}\right].$
In particular, $G_{\ell}(x)$ has no roots in the interval $[-1,1]$, which means that $R_{\ell}(s)$, and consequently $Q_{2\ell+1,2\ell-1}(s)$, are non-vanishing on the unit circle. Since $Q_{2\ell+1,2\ell-1}(s)$ is palindromic, Proposition 3.18 shows that exactly two of its roots lie inside ${\mathbb{D}}^{*}$. ∎
Remark 4.19.
As in the $m-n=1$ case in Remark 4.6, we can describe the limit behavior of the roots of $Q_{2\ell+1,2\ell-1}(s)$. Note that by (4.14), $s^{2}R_{\ell}(s)$ is a monic polynomial sharing the same roots as $Q_{2\ell+1,2\ell-1}(s)$. Now observe
$\lim_{\ell\to\infty}s^{2}R_{\ell}(s)$ $=1+6s+10s^{2}+6s^{3}+1=(s+1)^{2}(s^{2}+4s+1),$
which has roots at $\{-1,-1,-2+\sqrt{3},-2-\sqrt{3}\}$. Thus one interior root moves to the unit circle ($s=-1$) as $\ell\to\infty$, but the other interior root stays inside ${\mathbb{D}}^{*}$ and approaches $s=-2+\sqrt{3}$. Recall that this was the root accumulation point when $m-n=1$.
Corollary 4.20.
The Bergman kernel of ${\mathbb{H}}_{\frac{2\ell+1}{2\ell-1}}$ has zeros inside ${\mathbb{H}}_{\frac{2\ell+1}{2\ell-1}}\times{\mathbb{H}}_{\frac{2\ell+1}{2\ell-1}}$.
Proof.
Same proof as Corollary 4.7 with trivial modifications. ∎
4.3. Case $m-n\geq 3$
We now let $k\geq 3$, and briefly consider the setting of relatively prime pairs $m,n\in{\mathbb{Z}}^{+}$ with $m-n=k$.
Since $Q_{m,n}$ is a palindromic polynomial of degree $2k$, analysis becomes considerably harder as $k$ grows. However, there is ample evidence to suggest that every $Q_{m,n}$ is non-vanishing on the unit circle.
The authors have successfully carried out longer analogs of the argument given in Proposition 4.13 to prove the non-vanishing of $Q_{m,n}$ on the unit circle when $k=3,4,6$, which requires analysis of families of polynomials of degree $6,8,12$, respectively. Interestingly, the (still open) $k=5$ case is more difficult because (letting $\varphi$ denote Euler’s totient function; see [18]) $\varphi(5)=4$, while $\varphi(3)=2$, $\varphi(4)=2$, $\varphi(6)=2$. The higher the value of $\varphi(k)$, the more separate cases must be considered.
Though the problem for a general pair of relatively prime integers $m>n$ is still wide open, the authors have been able to prove that, for each $k\in{\mathbb{Z}}^{+}$, there are an infinite number of (explicitly given) pairs $m,n$ with $m-n=k$ such that $Q_{m,n}$ is non-vanishing on the unit circle. The argument requires detailed analysis of the polynomial coefficients considered in tandem with conditions which bound the number of real roots of cosine polynomials with real coefficients. While this material is beyond the scope of the present article, it is the subject of a new paper currently in preparation. In addition, the authors have used Mathematica to analyze an abundance of $Q_{m,n}$ polynomials not covered by these theoretical considerations and have never observed a single one with a root on the unit circle.
Conjecture 4.21.
Let $m>n$ be positive integers with $\gcd(m,n)=1$. The degree $2m-2n$ palindromic polynomial $Q_{m,n}$ is non-vanishing on the unit circle. Consequently, $m-n$ of its roots lie inside ${\mathbb{D}}^{*}$.
A proof of this conjecture would be somewhat remarkable as palindromic polynomials are generically expected to have roots on the unit circle. Indeed, an elegant and counter-intuitive result of Dunnage [10] says that, given a single randomly generated palindromic polynomial with real coefficients, the expected value of the percentage of its zeros lying on the unit circle is $\frac{1}{\sqrt{3}}$, or about 57.7%.
If Conjecture 4.21 can be proved, it would immediately imply that the Bergman kernel of ${\mathbb{H}}_{m/n}$ has zeros in ${\mathbb{H}}_{m/n}\times{\mathbb{H}}_{m/n}$ for every positive rational number $\frac{m}{n}\neq 1$.